OpenStudy (anonymous):

Which equations have a negative solution? Choose all answers that are correct. A. B. C. D. c d

OpenStudy (anonymous):

http://prntscr.com/8gz9fu

OpenStudy (madhu.mukherjee.946):

a &d

OpenStudy (jacobciezki):

That is what I was going to say cuz b is wrong

OpenStudy (jacobciezki):

so a & d

OpenStudy (texaschic101):

A. -w/2 = 11 --- multiply both sides by 2 -w = 22 w = -22 B. -1/6x = -3 x = -3/(-1/6) x = -3 * -6 x = 18 C. -4 = y/-10 -- multiply both sides by -10 40 = y D. -3/4z = 8 z = 8/(-3/4) z = 8 * -4/3 z = - 32/3

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

:)

OpenStudy (madhu.mukherjee.946):

yes

OpenStudy (jacobciezki):

Lol